it cost $10 to go to the movies. sarahs mom gave her $60 and told her she could invite friends to go see the hunger games when it came out. up to how many friends could sarah have invited?


1
Expert's answer
2021-09-17T03:48:53-0400

let the number of friends be x

therefore together with friend they will pay

"(1+X)\times10=60\\10+10X=60\\10X=60-10\\10X=50\\X=5"

Sarah can invite 5 friends